### v1

# Natural Language Generation from Visual Sequences: Challenges and Future Directions

Abstract
The ability to use natural language to talk about visual content is at the core of human intelligence and a crucial feature of any artificial intelligence system. Various studies have focused on generating text for single images. In contrast, comparatively little attention has been paid to exhaustively analyzing and advancing work on multiple-image vision-to-text settings. In this position paper, we claim that any task dealing with temporally ordered sequences of multiple images or frames is an instance of a broader, more general problem involving the understanding of intricate relationships between the visual content and the corresponding text. We comprehensively analyze five tasks that are instances of this problem and argue that they pose a common set of challenges and share similarities in terms of modeling and evaluation approaches. Based on the insights from these various aspects and stages of multi-image-to-text generation, we highlight several open questions and suggest future research directions. We believe that these directions can advance the understanding of complex phenomena in this domain and the development of better models.
1 Introduction
Over the years, research in natural language generation has demonstrated the importance of grounding language in the visual modality to improve understanding and reasoning capabilities of models Baroni (2016); Beinborn et al. (2018); Wang et al. (2024b). Earlier work on visually conditioned language generation primarily focused on single-image-to-text tasks such as image captioning and visual question answering. However, many practical real-world vision-to-language (V2L) applications in several domains such as surveillance and media content creation require understanding and reasoning across multiple temporally ordered images or video frames. With the increase in availability of video and image sequence data, various tasks have been proposed over the past few years to develop and evaluate models that can generate text grounded in multiple images or frames of videos. While some proposed multi-image-to-text tasks such as Video Captioning (VC) Yao et al. (2015) and Multi-image/Video Question Answering (VIDQA) Zeng et al. (2017); Bansal et al. (2020) are reminiscent of popular single-image-to-text settings, other tasks such as Change Captioning (CC) Jhamtani and Berg-Kirkpatrick (2018) are unique both in terms of their objectives and the type of input-output data.
Task Input Output
[…]
(a) Between consecutive images or video frames of input.
(b) Between consecutive sentences in ground-truth text.
Figure 1: Similarity scores obtained for the tasks along the visual and textual dimensions. We exclude Video Question Answering task from textual consistency analysis due to the lack of multi-sentence datasets.
Nevertheless, all multi-image-to-text tasks require models to reason along the temporal dimension of the visual input for generating the textual output. Therefore, in this position paper, we consider all multi-image-to-text tasks as instances of the broader problem of generating natural language output given a sequence of multiple temporally ordered images or video frames as input. Table 1 outlines a set of tasks that we consider as good representatives of this problem and presents the differences between the corresponding input and output data type. For all the tasks we provide examples in Table 2.
In summary, our main contributions are:
  • i.
    We quantitatively characterize each of the multi-image V2L tasks along two dimensions based on the complexities of their input-output data. We argue that the degree to which they vary along these dimensions is dependent on the corresponding task definitions and objectives.
  • ii.
    We highlight and discuss a common set of challenges relevant for multi-image-to-text generation such as accurate tracking and grounding of entities, ensuring coherence between output text segments, inter alia.
  • iii.
    We comprehensively review the evolution of modeling approaches, learning procedures, and evaluation protocols and provide a unified overview, which we believe will be useful for facilitating further advancements.
  • iv.
    Finally, we propose future research directions to improve the systems and methods used for tackling these tasks at different stages of the process.
2 Dimensions of Variation
While our focus is on the multi-image-to-text-problem, different tasks within this problem space may have different characteristics. Two relevant dimensions are the complexity of the visual input and of the textual output. For example, some of these tasks require models to generate succinct answers and descriptions, and others require generation of long-form textual narratives intended to also complement the visual information. These dimensions of variation across tasks typically tend to be crucial factors in making design choices with regard to developing model architectures and learning procedures. In terms of the visual input, depending on the objective of the task, images within the input sequence of each data sample could be comparable to each other or vary drastically to the point of being completely heterogeneous. For instance, in the CC task, where the goal is to localize and describe changes between a pair of images obtained from real-time surveillance cameras or large-scale remote-sensing snapshots, we hypothesize that the similarity between input images would be generally high. On the other hand, in tasks such as Visual Storytelling (VST) Huang et al. (2016), in which the input sequences typically depict an overarching narrative, we hypothesize low similarity between consecutive images within each data sample (<visual sequence, text> pair). Regarding textual output, we similarly posit that the consistency of consecutive sentences within each data sample could be high or low depending on the corresponding task objective.
To preliminarily test our intuitions, we quantitatively analyze a few datasets corresponding to each of these tasks. Specifically, for each data sample, we compute visual similarity scores between CLIP Radford et al. (2021) visual encoder embeddings of consecutive images in the input sequence and report the average score. In the same manner, we compute textual consistency scores between CLIP text encoder embeddings of consecutive sentences in the corresponding ground-truth text of each data sample. For this study, we randomly select 100 instances per task from five datasets—Spot-the-diff Jhamtani and Berg-Kirkpatrick (2018) for CC, VIST Huang et al. (2016) for VST, Charades Sigurdsson et al. (2016) for VC, MSVD-QA Xu et al. (2017) for VIDQA, and MAD-v1 Soldan et al. (2022) for Movie Auto Audio Description (Auto AD) Han et al. (2023b) (MAAD). [^1] Further details are provided in Appendix A. Figure 1 shows the distributions of similarity scores obtained for each of the tasks along the visual and textual dimensions. In terms of visual similarity, we observe that CC and VST indeed obtain maximum and minimum scores respectively, with other tasks ranging in between, confirming our intuitions. In terms of textual consistency, the differentiation is less evident. We observe that for the MAAD and VC tasks, consecutive sentences in the ground-truth text are relatively less consistent with each other. Using the average similarity scores across all data samples, we categorize the five tasks by placing them at different positions in the shared space between textual consistency and visual similarity (see Figure 2). Besides the two axes considered for this analysis, we note that tasks could also be compared along various other dimensions. We posit that this kind of analysis would serve as a meaningful guide for making modeling and evaluation decisions both for current and for novel future tasks in the multi-image-to-text landscape.
Figure 2: Tasks positioned in the visual-textual shared space using the similarity scores obtained for corresponding datasets. Interpretation: a task in the top right corner of the plot is both inconsistent at the textual level and dissimilar at the visual level.
3 Common Challenges
Multi-image-to-text tasks present a unique set of challenges. Some of these challenges are task-specific, but most of them are common across the five tasks. In this section, we describe these challenges and discuss how they instantiate for each of the tasks.
Entity tracking.
Identifying and tracking entities is an important requisite for accurately interpreting actions and relationships between them. This has been underlined by several works in both language understanding Paperno et al. (2016); Kim and Schuster (2023) and computer vision domains Narayan et al. (2017); Luo et al. (2021). In multi-image V2L tasks, the availability of input signals from two modalities makes the aspect of disambiguating entities more challenging. For instance, in the MAAD or VST tasks where the visual input is heterogeneous, entities in the input image sequences/video clips tend to ‘disappear’ in some of the images/frames at the intermediate temporal positions, while still being actively referenced in the textual input at the corresponding positions. Surikuchi et al. (2023) have denoted such cases as being temporally misaligned. To track entities accurately under such temporal misalignment, it is necessary to not only learn causal relations between people and objects in each of the modalities at corresponding positions, but also to obtain a cross-modal cross-temporal representation of all the relationships relevant to the scene/narrative. For the CC, VC, or VIDQA tasks, in which the input images/video frames are typically similar to each other, it is crucial to differentiate meaningful semantic entities and their changes from various distractions. While in CC, viewpoint changes or illuminations are considered as distractors and discarded, in the VIDQA task, entities relevant to answering the question need to be differentiated from others for accurate tracking. Effective tracking of entities would therefore require accounting for changes in appearance (including disappearance), capturing interactions, and correctly identifying occlusions.
Visual grounding.
Humans acquire language understanding through perception and interaction with the environment Barsalou (2008); Iverson and Goldin-Meadow (2005) and consequently this enables them to seamlessly ground language in visual data. Over the years, a great deal of work has been proposed to adapt the architecture and learning process of vision-language models for acquiring visual grounding Suglia et al. (2024). However, there are still significant challenges for achieving human-levels of grounding using computational models and this becomes more apparent when looked at from the perspective of multi-image-to-text tasks. For instance, in the VST or MAAD tasks, since language is typically ‘inconsistent’ (see Figure 2), it is also inadvertently under-specified semantically Pezzelle (2023) (e.g., ‘The boy on the bridge was waving to the tourists near the waterfall. A photographer over there clicks…’). Stories also tend to contain many abstract adverbs such as ‘often’ or ‘today’ and it has been shown that vision-language models struggle to disambiguate such under-specified text and accurately map phrases to regions in the image sequences/videos. Moreover, the amount of language informativeness—degree of information required for identifying the correct object Coppock et al. (2020)—could be inadequate in tasks such as VIDQA, particularly with the presence of confounding entities in various frames (e.g., input video of a football match and a question: ‘What is the color of the card the referee is holding?’). Also, when grounding objects, it has been shown that models often struggle to reliably capture spatial relationships Kamath et al. (2023). To summarize, beyond being merely descriptive, language in some multi-image-to-text tasks could also be complementary to the data in images/videos, making visual grounding challenging without access to relevant additional external knowledge.
Knowledge integration.
For some multi-image-to-text tasks, models would be required to utilize additional information beyond what is available in the input data. In VC or VIDQA tasks pertaining to certain domains such as news, the input video might not contain all the aspects needed to correctly describe its contents or answer questions about it Whitehead et al. (2018); Jin et al. (2023). To address this, various approaches often rely on using large pre-trained general purpose models or external knowledge bases such as ConceptNet Speer et al. (2017) to retrieve both factual and commonsense information. This method is commonly referred as retrieval-augmented generation (RAG) Lewis et al. (2020). Besides being sources for missing information, external knowledge bases are also leveraged for enriching the generated text with social or cultural contexts. For instance, in VST, some approaches use recognized visual objects in the input to retrieve concepts from external knowledge graphs for generating more engaging and figurative stories (e.g., concept of ‘graduation ceremony’ following the detection of an ‘academic gown’ object).
From knowledge selection/retrieval stage to accurately representing and utilizing it during text generation, the process of integrating external knowledge has various challenges. Robust retrieval systems are required which can holistically extract the essence of image sequences/video frames, including the various entities and their interrelationships. Typically, the retrieved knowledge is concatenated with input representations which are then used for generating text either through fine-tuning Yang et al. (2024) or by prompting general-purpose VLMs Bhattacharyya et al. (2023). However, this approach might lead to models either over or under utilizing the retrieved knowledge potentially leading to incoherent text Gao et al. (2023). To address this, we argue that fusion mechanisms which can effectively balance information from representations of both input data and the retrieved knowledge need to be developed. Furthermore, retrieving relevant knowledge from increasingly large databases could be computationally expensive, especially in the multi-image scenario. Ways to optimize retrieval components for improving efficiency is an active research area.
Textual coherence.
Coherence is the property of text that refers to the ordering of its constituents (words/sentences) and the way in which they relate to each other Althaus et al. (2004). Coherent text should have a consistent logical structure in which the events, interactions, and relationships between various elements are ordered in a meaningful way. It is an important aspect of discourse and has been studied extensively in neural language generation Pishdad et al. (2020). For several multi-image-to-text tasks, particularly the ones in which the output tends to be less ‘consistent’ along the text axis in Figure 2, it is challenging to ensure that multiple sentences in the generated output are locally coherent. In MAAD and VST tasks, where there are multiple characters and various interactions developing across the sequence of images/video frames, it is often difficult for models to balance between selecting the visual information and representing it in a cohesive/connected language Lei et al. (2020). This challenge is more apparent in the VST task, in which models are expected to keep track of multiple things such as emotional arcs or motivations of the characters and the overarching narrative Surikuchi et al. (2024). There is increasing work in unimodal text-only storytelling suggesting how using concepts of narratology Piper et al. (2021); Antoniak et al. (2024) such as Genette (1980)’s triangle can potentially aid models in generating stories with engaging and coherent structures. However, it still unclear how these theories can be applied to multimodal scenarios where the generated text needs to be consistent with image sequences/videos.
Theory of mind.
Theory of Mind (ToM), which is considered the basis of human social cognition Premack and Woodruff (1978), is described as the ability to understand and make inferences about the mental states (e.g., beliefs, intentions, and desires) of other people or living beings. In the context of multi-image V2L tasks, ToM refers to the ability of models to go beyond merely recognizing objects/actions and to reason about the mental states of entities depicted in the image sequences/videos. Although most of the tasks we consider in this work, besides VIDQA Mao et al. (2024), do not explicitly require ToM abilities, the skill is still relevant for all the tasks and closely connected to the other challenges and abilities discussed so far. For instance, in tasks such as VST, to causally connect heterogeneous images in the input sequence, models need to be equipped with different reasoning abilities pertaining to emotions and social perceptions/intentions. This enables generation of stories that reflect actions and mental states of the characters beyond literal interpretation of the visual data.
Recently, several ToM benchmarks have been proposed to assess general-purpose VLMs Gao et al. (2024); Chen et al. (2024b) along different aspects such as temporal localization of emotions, intentionality-understanding, and perspective-taking. However, these studies find that only models that are fine-tuned on curated ToM datasets exhibit any reasoning abilities, albeit not aligning with the well-established ToM theories/frameworks explaining human social cognition. Such curated data is scarcely available and it is unclear what alternative architectures or training objectives would enable models to obtain the ToM abilities required for multi-image V2L tasks.
4 Models Architectures
Figure 3: Outline of the architecture common across modeling approaches for multi-image-to-text tasks.
Modeling approaches to multi-image-to-text tasks have evolved over time from being recurrent neural network (RNN)-based Hochreiter and Schmidhuber (1997) to being transformer-based Vaswani et al. (2017). More recent models directly leverage pre-trained large (vision)-language models (LLMs/VLMs), often in a zero-shot manner. In this section, we discuss this evolution and summarize the various state-of-the-art model architectures proposed for the five multi-image-to-text tasks. Architectures proposed for these tasks primarily comprise three modules—a vision encoder, a language decoder, and an intermediate module (typically referred as the projector/adapter) for adapting visual information into contextualized representations for text generation. We describe the functionality of these modules and review the design principles common across all the tasks in the proposed approaches. Furthermore, we also discuss how off-the-shelf pre-trained VLMs are currently being used to handle various multi-image-to-text tasks. Table 3 outlines a summary of the models and details related to the selection procedure are provided in Appendix B.
4.1 Vision Encoder
The primary purpose of a vision encoder in vision-to-language tasks is to extract information from the input visual sequence and to optimally encode it into a contextual representation that guides language generation. To achieve this, encoders in the proposed models follow multiple steps, some of which are common across the five multi-image-to-text tasks. First, a pre-trained vision model is utilized for extracting feature representations of the raw input sequences of images/video frames. Earlier approaches used convolutional neural network (CNN)-based vision models such as ResNet He et al. (2016) or R3D Tran et al. (2018) that are primarily pre-trained on the object detection task using large amounts of image/video data. Most of the recent models across the tasks use transformer-based vision models pre-trained for various image-only and image-text alignment objectives, e.g., CLIP-ViT-L Radford et al. (2021).
We note that besides the primary input sequence of images/video frames, models proposed for some of the tasks, e.g., MAAD, utilize additional input data such as close-ups of characters in the movie clips (exemplars) Han et al. (2023a). Furthermore, the TAPM Yu et al. (2021) model proposed for the VST task utilizes FasterRCNN Ren et al. (2015) for extracting such local character/object-level features alongside the global image-level features from ResNet. Following the extraction of visual features using pre-trained vision models, most vision encoders comprise an internal sequence-encoder component for learning relationships and dependencies between the individual image/frame-level features at different temporal positions. Some models implement this step either using RNNs or a transformer network with multi-head self-attention for learning temporal relationships and position-encoding mechanism for tracking the order of entities/events in the sequence.
Beyond these steps that are common across the tasks, vision-encoders may also contain additional task-specific steps for capturing the visual information in a way that suits the task’s objective better. For instance, the ViLA model Wang et al. (2025) for the VIDQA task utilizes a learnable Frame-Sampler component to efficiently select a small subset of frames that are most likely to contain the relevant information needed to answer the question. Another example with a task-specific step is the MSCM+BART model Hsu et al. (2020) for VST, in which the initial set of image objects/‘concepts’ are expanded using an external knowledge graph for generating diverse and informative stories. Despite these task-specific steps, we found that the vision encoder module in the architectures proposed for the various multi-image-to-text tasks share a common set of components that are broadly outlined in Figure 3.
4.2 The Vision-to-Language Bridge
Some V2L model architectures utilize an intermediate module that bridges the input and output modalities for effectively conditioning the text generation on the extracted visual features. Different models operationalize this module with different degrees of complexity. Earlier approaches for several multi-image-to-text tasks condition the text generation process by directly fusing vision encoder outputs with the language decoder input Kim et al. (2018). Some architectures employ cross-attention mechanisms to focus on the relevant parts of the visual features at various temporal positions during decoding Yao et al. (2015). However, approaches that adopt pre-trained models—e.g., CLIP-ViT-L Radford et al. (2021) as the visual model—tend to employ learnable intermediate layers for aligning and converting outputs of the vision encoder into a format that the language decoder can understand.
In some of the proposed models, this intermediate module is a single linear layer that transforms the visual features into a common shared space, which can be used by the language decoder Ko et al. (2023); Liu et al. (2023). In other models, advanced transformer-based projectors such as a Q-Former Li et al. (2023c) are used for their ability to leverage cross-modal interactions effectively Han et al. (2024). In essence, Q-Former uses dynamic query vectors that are pre-trained to attend to both visual and textual representations, enhancing its ability to generalize and perform well (relative to a single linear layer) across different tasks. Besides these popular methods for adapting multimodal information, some approaches make use of graph neural networks for capturing relationships between objects in the images at different temporal positions and words in the corresponding sentences of the text Zhang and Peng (2019). While there is no definitive way to design this intermediate module, recent work has compared the two approaches, i.e., using cross-attention between modalities or using a multimodal projector for transforming vision encoder features into the language space, and found that the latter leads to a stable/improved performance of models Laurençon et al. (2024b).
4.3 Language Decoder
After encoding and adapting the visual information, models employ a language decoder component for text generation. The decoder can either be learned from scratch or consist of a pre-trained language model with additional trainable task-specific layers. Figure 3 summarizes the different ways in which this step is operationalized across tasks in the proposed architectures. Earlier models learn an RNN by initializing it with the visual context embedding from the previous steps Yao et al. (2015); Kim et al. (2018). The decoder then typically follows a ‘teacher forcing’ strategy during training to generate one word at a time autoregressively.
Subsequent models have replaced RNNs with the transformer architecture owing to its computation scalability and efficiency in handling long context-windows. Besides the initial word embedding layer and the position encoding step (for maintaining information about the input sequence token order), a transformer decoder is typically made up of multiple identical blocks. Each block comprises a multi-head self-attention layer for modeling intra-sentence relations (between the words) and a multi-head cross-attention layer for learning relationships between representations of each word and the outputs of the visual encoder/projector. For instance, in the CC task, this refers to conditioning each word in the caption on vision encoder outputs (denoted as ‘difference-representations’).
Instead of training the decoder from scratch, some approaches use language models such as GPT-2 Radford et al. (2019) and LLAMA 2 Touvron et al. (2023), which are pre-trained on several text-only tasks such as question-answering and text classification/completion. The pre-trained language models are either used directly for generation by freezing their parameters Han et al. (2023b, a, 2024), or by inserting and fine-tuning additional adaptive layers on top of them for ensuring relevance of the generated text to the downstream task of interest Yu et al. (2021). We also note that some models incorporate information from external knowledge bases/graphs into the decoder module to improve coherence and factuality of the generated text, e.g., TextKG Gu et al. (2023b) for the VC task and KG Story Hsu et al. (2020) for the VST task.
4.4 Off-the-shelf Pre-trained VLMs
The standard model architecture we have discussed so far is also present in more powerful general-purpose foundation VLMs (pre-trained on several tasks using large amounts of data), which can be used directly for multi-image-to-text tasks. Their pre-training process typically happens in two stages—self-supervised alignment training and visual instruction tuning. During the first stage, only the parameters of the intermediate module connecting both unimodal backbones are updated (commonly using paired image-text data) utilizing a contrastive training objective.
In the second stage, models are instruction-tuned using multi-turn conversations obtained for visual data either through crowd-sourcing or by leveraging tools such as GPT-4 OpenAI (2023). Contrary to task-specific modeling approaches, these pre-trained VLMs are simply prompted (typically in a zero-shot manner) using visual tokens accompanied by task-specific instructions. Some of the pre-trained VLMs that are used off-the-shelf for the multi-image-to-text tasks include: ViLA for VIDQA, mPLUG-2 for VC, VideoLLAMA for MAAD, and LLaVA-NeXT for VST Wang et al. (2025); Xu et al. (2023); Xie et al. (2024); Surikuchi et al. (2024).
5 Evaluation
Given all the similarities described above, it is not surprising that all multi-image-to-text tasks are also evaluated leveraging similar methods. These methods range from using traditional n-gram matching metrics to obtaining human judgments and ratings to, more recently, using off-the-shelf pre-trained VLMs assessing the generated output. We broadly classify these evaluation methods into two main categories—automatic and human evaluation. In the following subsections, we discuss the several quantitative metrics and benchmarks widely used for each of the tasks, along with the rationales for relying on them.
5.1 Automatic Evaluation
To computationally assess the quality of model-generated text along different aspects, several automatic metrics have been proposed. While some metrics rely on answers/text provided by human annotators, others are reference-free and assess model outputs independent of the ground-truth data. Besides computational metrics, the community has also relied on benchmark datasets designed to reveal various general capabilities of models.
Reference-based metrics.
The five multi-image-to-text tasks we examined primarily assessed model-generated candidate text by comparing it to corresponding human-written references. Specifically, traditional metrics that were originally designed for evaluating machine translation and text summarization tasks—BLEU Papineni et al. (2002), METEOR Banerjee and Lavie (2005), and ROUGE Lin (2004)—are used to measure precision and recall of overlapping $\mathit{n}$-grams between the candidate and reference text. Usually, metrics such as CIDEr Vedantam et al. (2015) and SPICE Anderson et al. (2016) that have been specifically developed for the evaluation of image and video captioning are also used in conjunction with the above three metrics.
All the above-mentioned metrics rely on direct raw text comparisons of ground-truth references and model outputs. As this ground-truth data might not be always available, embedding-level reference-based evaluation metrics such as WMD Kusner et al. (2015), BERTScore Zhang* et al. (2020), and ViLBERTScore Lee et al. (2020) have been proposed. Recent work for VC, VIDQA, and MAAD tasks have incorporated these metrics to measure the similarity between candidate and reference embeddings, obtained by projecting corresponding text into a common pre-trained semantic space Lin et al. (2022); Han et al. (2023a, 2024).
Reference-free metrics.
Comparing model-generated text to ground-truth references, typically provided by crowd-workers or scraped from the internet has various limitations. Most reference-based metrics do not account for the visual modality upon which the generated text is conditioned. Furthermore, as shown in Figure 2, text in the VC, MAAD, and VST tasks often complements the visual input by encompassing various abstract/creative concepts, and is not merely descriptive. This makes reference-based metrics generally inappropriate for accurately evaluating multi-image-to-text tasks.
For the reasons detailed above, various reference-free metrics such as MAUVE Pillutla et al. (2021) and UNION Guan and Huang (2020) have been proposed for unimodal open-ended text generation tasks. However, for the visually-conditioned text generation tasks, this adoption/shift is still relatively recent, with a persistent emphasis on reference-based $\mathit{n}$-gram metrics to date. That said, various reference-free metrics have been recently proposed to assess different aspects of evaluation that are important for several tasks. For instance, metrics such as CLIPScore Hessel et al. (2021) and GROOViST Surikuchi et al. (2023) have been developed for evaluating visual grounding—the degree of alignment between the generated text and the visual input—in VC and VST tasks. Similarly, the RoViST Wang et al. (2022) suite of metrics has been proposed to assess coherence, the extent of repetition, and visual grounding of the generated text. We note that there are also metrics such as CRITIC Han et al. (2024) and CM (character matching) Hong et al. (2023) designed to evaluate task-specific aspects such as the accuracy of referencing to characters in the model outputs. Besides the above-mentioned metrics (tailored to measure specific features of the generated content), there is also an increasing adoption and reliance on using pre-trained LLMs and VLMs as judges Li et al. (2024a). Essentially, these pre-trained general-purpose models are prompted to score or rate a model-generated response along any of the evaluation dimensions of interest, such as fluency or relevance (e.g., ‘How fluent is the generated text on a scale of 1 to 5?’). However, the effectiveness and reliability of this approach is still debated Bavaresco et al. (2024).
Benchmarks.
With the increase in scale, data, and extensive multi-step training processes it is difficult to fully understand the capabilities of models based only on their performance on held-out test splits of task-specific datasets. To address this limitation, numerous benchmark datasets have been proposed that focus on evaluating more fine-grained abilities of models. Some of the popular benchmarks proposed to test models trained for multi-image scenarios include: NLVR2 Suhr et al. (2019) which focuses on models’ ability to understand the visual compositionality given a pair of images and corresponding textual description; ViLMA Kesen et al. (2024) and MVBench Li et al. (2024c) which focus on testing models’ spatio-temporal reasoning capabilities (e.g., counting actions across frames of a video); Mementos Wang et al. (2024a) which studies object and behavioral hallucinations, and their interconnectedness.
Despite continued progress to improve and update existing benchmark datasets to cover various edge cases, evaluation using benchmarks is not without its limitations. For instance, some benchmarks are constructed using data from test/validation splits of existing popular datasets in the community, leading to a potential contamination problem Laurençon et al. (2024a). Moreover, recent modeling approaches typically tend to incorporate most existing benchmark datasets into their fine-tuning process to ensure the stability and generalization of models in real-world applications. When such approaches abstain from disclosing the data used for training the models, it generally undermines the process of automatic evaluation using benchmarks for comparing models against each other.
5.2 Human Evaluation
Given the current state of automatic evaluation, some multi-image-to-text tasks such as VC and VST rely on human evaluation to accurately determine the quality of the model-generated text. This process involves recruiting online crowd-workers who are native or proficient speakers of the target language. Depending on the type of data or variation of the task, annotators with expertise and familiarity with terminology relevant to the corresponding domain (e.g., medical or sports videos) might be preferred.
Participants of the evaluation study are provided with a set of task-specific rubrics/instructions along with representative examples required for judging the model outputs. They are asked to assess the overall quality of model-generated outputs either independently (per sample) Surikuchi et al. (2023) or relative to outputs from other models Wang et al. (2022). Alternatively, evaluators might be required to provide scores/ratings for various criteria ranging from broad (e.g., text conciseness, fluency, grammatical correctness) to specific (e.g., factuality, hallucinations, expressiveness). The obtained scores are usually compared pairwise to rank models appropriately.
Some pre-trained VLM frameworks such as LLaVA-RLHF Sun et al. (2024) leverage this qualitative feedback to optimize model parameters for learning to generate human-preferred text. Although human evaluation is still indispensable for several tasks, it is also expensive, time-consuming, and challenging. Defining clear evaluation protocols for ensuring the reliability and quality of human judgments is an active research area Kasai et al. (2022); Ruan et al. (2024).
6 Discussion
As discussed above, the problem of generating text from a sequence of temporally ordered images or frames is a challenging one, and relevant to several downstream tasks and applications. Here, we reflect on some crucial aspects and outline various prospective research directions (RDs) and takeaways.
RD 1: Towards more naturalistic scenarios.
Many of the multi-image-to-text tasks we consider in this work have real-world applications. For instance, solutions to the CC task can be used for assisted surveillance and for tracking changes in digital media assets Jhamtani and Berg-Kirkpatrick (2018). In the MAAD task, models are required to generate descriptions that complement information in the original audio dialog/soundtrack, for improved accessibility to visually impaired users and for enhancing the visual experience of sighted users Han et al. (2023a).
However, many day-to-day human-centered scenarios involve personalizing to various contexts or situations. We argue that existing multi-image-to-text tasks in their definitions and settings do not fully reflect this aspect. Tailoring model-generated descriptions/narrations to the perspective of end-users requires task settings in which models and humans can interact iteratively. Such settings would enable incorporation of human expectations and communicative contexts which typically tend to be dynamic in real-world applications. To this end, we advocate for variations of existing tasks where models can learn to contextualize and reason through interactions with other agents (humans or other models) for generating stories, descriptions, or answers. Furthermore, we also advocate for exploration of controlled task settings in which models are expected to generate text adhering to a specific style Yang and Jin (2023) or point-of-view.
RD 2: Are general-purpose VLMs all we need?
As discussed in Section 4.4, VLMs that are trained on various general-purpose datasets are increasingly being used for multi-image-to-text tasks through prompting. Powerful open-source models such as Molmo Deitke et al. (2024), and models optimized for multi-image scenarios such as Mantis Jiang et al. (2024) are becoming increasingly available, suggesting that the trend of adopting them off-the-shelf for solving many V2L tasks is widespread. General-purpose VLMs learn abundant information through multitask pre-training and have a modular design, making them suitable for many downstream tasks. Their modularity also enables for seamless adaptation of VLMs to various novel domains (e.g., medical science) by updating only a small fraction of their parameters Li et al. (2023a).
Despite the promising generalization of VLMs to certain tasks and domains, they have also been shown to be sensitive to prompts Liu et al. (2024); Schlarmann and Hein (2023) and biased towards the textual modality Rahmanzadehgervi et al. (2024). To address these problems, recent work proposes various prompt engineering techniques to facilitate inference-time adaptation of prompts to make them more suitable for the specific task of interest Ma et al. (2023); Gu et al. (2023a). On the other hand, task-specific model architectures consist of components designed to effectively address specialized aspects of the tasks, e.g., computing a difference representation of the input image pair in CC. We advocate for modular modeling approaches that bring together efficient task-specific components and combine them with the powerful foundational VLMs. Furthermore, we argue that using graph-based architectures and memory-based modules would result in improved tracking of entity positions/relationships and enable models to assign saliency to memorable events in tasks like VST or MAAD.
RD 3: Improving and rethinking evaluation.
In Section 5, we discussed the various approaches for evaluating model outputs in multi-image-to-text tasks. While human evaluation is impractical for conducting large-scale assessments, existing automatic evaluation metrics are limited in terms of fully capturing the abilities of models. Increasingly various benchmarking datasets are being proposed to assess models along different axes important for grounding language in the visual input Li et al. (2024b). However, many benchmarks often suffer from the problem of visual content irrelevance, which refers to models performing well on the benchmark datasets by primarily relying only on the language modality Chen et al. (2024a). Furthermore, data leakage and contamination problems (see section 5.1) also hinder fair and accurate testing of model’s skills using benchmarks.
While it is important to continue directing research efforts towards developing more extensive multi-image benchmarks such as ReMI Kazemi et al. (2024), we argue that the purpose of evaluation is to also provide insights that can be directly leveraged for improving model architectures and learning procedures. For single-image-to-text tasks, recent works have adapted interpretability methods that focus on understanding the behavior and internal representations of models Neo et al. (2024); Yu and Ananiadou (2024). These methods can complement traditional evaluation techniques for enabling intra- and inter-model comparisons of behaviors and mechanisms for obtaining a holistic understanding. We strongly advocate for work that adapts various categories of interpretability methods for multi-image-to-text scenarios.
7 Conclusion
In this position paper, we focused on the problem of multi-image V2L generation and connected the various tasks that are typically considered separate by the research community. We proposed a method for quantitatively exploring the current landscape of this problem, using which we uncovered relationships between task objectives and their corresponding datasets. Despite having different characteristics in terms of the objectives or the input-output data, we argued that all the tasks present a common set of challenges for developing models and for assessing their generated outputs. To understand the progress made over the years with regard to multi-image-to-text generation, we extensively reviewed the different modeling approaches and evaluation protocols pertaining to each of the tasks, and discussed them in a comprehensive and unified manner. As the problem of generating text conditioned on sequences of multiple temporally ordered images or video frames has various real-world applications, we underline the challenges and propose several concrete research directions informed by insights from linguistics, cognitive sciences, and natural language processing (NLP) aimed towards facilitating further advancements. We argue that leveraging these insights could also help the development of better VLMs, which are currently not immune from some of the highlighted limitations.
Acknowledgments
We are immensely grateful to our colleagues at the Dialogue Modelling Group (DMG) for their invaluable suggestions at different stages of this work. We thank Alberto Testoni and Anna Bavaresco for their insightful feedback. Raquel Fernández was supported by the European Research Council (ERC Consolidator Grant DREAM 819455).
Appendix A Quantitative Analysis of Tasks
For computing similarity scores, we used the image and text encoders of the ViT-B/32 version of the CLIP model. Obtained visual similarity and textual consistency scores are subtracted from $1.0$ for representing the tasks on the shared space in Figure 2. For the Multi-image/Video Question Answering task, we set textual consistency value to $1.0$ due to the unavailability of datasets having multiple-sentences in the textual outputs.
Appendix B Models
Table 3 outlines the models we reviewed for each of the multi-image-to-text tasks in this work. To reflect the evolution of models, we considered both RNN-based approaches and the more contemporary Transformer-based architectures that obtained state-of-the-art results.
{tblr}
colspec = X[c,h]X[c]X[c]X[c], stretch = 0, rowsep = 5pt, hlines = xkcdBlack, 0.5pt, vlines = xkcdBlack, 0.25pt
& Change Captioning, dataset: Spot-the-diff
Caption: The blue truck is no longer there. A car is approaching the parking lot from the right.
Video Question Answering, dataset: IntentQA Li et al. (2023b), Question: Why did the man point to the screen when talking to the child?
Answer: Draw child’s attention.
Video Captioning, dataset: Charades
Caption: The person is sitting at the dining room table wrapped in a blanket. The person is eating cereal and drinking orange juice.
Movie Auto Audio Description, dataset: MAD-v1
Story: Sully adjusts his seat harness. A male passenger looks up from his magazine … Sully sticks out an arm as the jet bellies down onto the river.
Visual Storytelling, dataset: VIST
Story: A discus got stuck up on the roof. Why not try getting it down with a soccer ball? … It didn’t work so we tried … are all stuck on the roof.
Table 2: Examples for each of the multi-image-to-text tasks we consider in this work.
Model Vision Encoder Projector Language Decoder
Change Captioning
CARD Tu et al. (2024) ResNet, Transformer NA Transformer
SCORER+CBR Tu et al. (2023b) ResNet, MH(S/X)A NA Transformer
VARD-Trans Tu et al. (2023a) ResNet, Linear(s) NA Transformer
DUDA Park et al. (2019) ResNet, RNN NA RNN
Multi-image/Video Question Answering
ViLA Wang et al. (2025) ViT, Transformer Q-Former Flan-T5 XL
LLaMA-VQA Ko et al. (2023) CLIP-ViT-L Linear LLAMA
SeViLA Yu et al. (2023) ViT Q-Former Flan-T5 XL
FrozenBiLM Yang et al. (2022) CLIP-ViT-L Linear DeBERTa-V2-XL
Video Captioning
Vid2Seq Yang et al. (2023) CLIP ViT-L, Transformer NA T5-base
TextKG Gu et al. (2023b) Transformer NA Transformer
VTAR Shi et al. (2023) InceptionResNetV2, C3D Transformer Transformer
ENC-DEC Yao et al. (2015) 3DCNN Attention RNN
Movie Auto Audio Description
MM-Narrator Zhang et al. (2024) CLIP-ViT-L NA GPT-4
AutoAD-III Han et al. (2024) EVA-CLIP Q-Former LLAMA 2
AutoAD-II Han et al. (2023a) CLIP-ViT-L NA GPT-2, MHXA
AutoAD Han et al. (2023b) CLIP-ViT-L Transformer GPT-2
Visual Storytelling
MCSM+BART Chen et al. (2021) ResNet, RNN NA BART
TAPM Yu et al. (2021) ResNet, FRCNN NA GPT-2
KG Story Hsu et al. (2020) FRCNN NA Transformer
GLAC Net Kim et al. (2018) ResNet, RNN NA RNN
Table 3: A selection of recent models proposed for each of the tasks considered in the paper. All the models reported are end-to-end and task-specific, i.e., trained or fine-tuned for the task. For each model, we report the underlying vision encoder and language decoder, as well as the projector module, when applicable (NA: Not Applicable). Components with have been trained from scratch using only the datasets available for the corresponding task.